A spherically symmetric charge distribution produces the electric field E⃗  =( 4600 r2)r^N/C, where r is in m.

What is the electric field strength at r = 15.0 cm ?

What is the electric flux through a 30.0-cm-diameter spherical surface that is concentric with the charge distribution?

How much charge is inside this 30.0-cm-diameter spherical surface?
Answer #1

Ans:

  1. E = 4600r2 N/C

E = 4600(0.15)2 = 103.5 N/C

  1. 30 cm = Diameter

15 cm = Radius

φ = EAsphere

            = (103.5)(4 * Pi * (0.15)2) = 29.249 Nm2/C

  1. For charge

φ = q/ϵ0

q = φ * ϵ0 = (29.249)(8.85 * 10-12) = 2.58 * 10-10 C